OpenStudy (daniellelovee):

An airplane is flying 22 degrees north of east at a speed of 810 km/h. How fast is it moving to the north?

OpenStudy (daniellelovee):

I got 750 is that right?

OpenStudy (daniellelovee):

@Directrix

Directrix (directrix):

>I got 750 is that right? Where is the work for this. I don't know Physics well at all.

Directrix (directrix):

From Phone-A-Friend @Douglas The northern component (Vn) of the velocity is: Vn = V{sin(Θ)} Where V is the magnitude of the Velocity (a.k.a the speed) and Θ is the angle North of East. Vn = 810{sin(22º)} = ?

Directrix (directrix):

@mamabama disagrees and offers this: Difficult to explain without a triangle but here: The north component of the planes velocity is 810sin(22) = 303.4 The east component is 810cos(22) = 751.0 So the plane is moving north at 303 km/h

OpenStudy (daniellelovee):

thank you so much and im sorry but my computer froze for like half an hour -.-'

Directrix (directrix):

Don't assume that either of these is correct without checking.

OpenStudy (daniellelovee):

yeah but I still believe it has to be 750 :/

OpenStudy (anonymous):

\[\huge v_\text{North}=|v| \sin (\theta)\] By triangular orthogonal decomposition.

OpenStudy (anonymous):

@Directrix Both answers you posted are the same

OpenStudy (daniellelovee):

not really one is 750 and the other one is 810

OpenStudy (anonymous):

@Daniellelovee No His first post: `Vn = 810{sin(22º)} = ?` His second post: `The north component of the planes velocity is 810sin(22) = 303.4. So the plane is moving north at 303 km/h` They are the same since 810sin(22)=303.4

OpenStudy (shamim):

East component of the plane velocity is=810*cos22=?

OpenStudy (shamim):

And north component of the plane velocity is=810*sin22=?

OpenStudy (shamim):

U hv to know perpendicular vector resolving